Circular No. 6066

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET MACHHOLZ 2 (1994o) H. Lüthen, Hamburg, reports the discovery by Michael Jäger of a sec-

ond comet in the field of comet 1994o, located ∼ 48′ to the northeast of 1994o and appearing to have the same motion as comet 1994o. Additional observations have been reported in response to requests for confirmation from the Bureau:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Aug. 28.03611 7^{h}19.^{m}2  +47^{o}53′  11  Jäger
30.05970 7 31 53.^{s}75 +45 24 44′′.1  Varady
30.09678 7 32 06.78 +45 22 04.5 12.1  _{”}
30.09861 7 32.2 +45 21 Jäger
30.46925 7 34 15.52 +44 55 07.2  Balam
30.51247 7 34 30.46 +44 51 59.3  _{”}

M. Jäger (Vienna, Austria). 0.20-m Schmidt camera + 2415 Tech Pan film.

Diffuse coma 4′–5′ in diameter, with little or no condensation. M. Varady and P. Pravec (Ondřejov). 0.65-m f/3.6 reflector + CCD. Comet

has moderate central condensation; m_{2} = 16.3; coma diameter 2′.2, with a faint halo of diameter 3′.5.

D. D. Balam (Climenhaga Observatory, Victoria). 0.5-m reflector + CCD. Comet very diffuse, with no central condensation; difficult to measure.

The following ephemeris is from the 1994o orbit on IAUC 6059, with T and ω changed to 1994 Sept. 18.208 and 147^{o}.497:

1994 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Aug. 26 7^{h}04.^{m}48 +50^{o}29.^{′}1 0.408 0.854 56.^{o}2 100.^{o}4 11.4
28 7 18.97 +47 57.2 0.426 0.838 54.6 101.0 11.4
30 7 31.52 +45 29.3 0.444 0.824 53.2 101.3 11.4
Sept. 1 7 42.56 +43 06.5 0.464 0.811 52.0 101.2 11.4
3 7 52.41 +40 49.2 0.484 0.798 51.0 100.9 11.4
5 8 01.32 +38 37.5 0.505 0.788 50.2 100.2 11.5
7 8 09.49 +36 31.5 0.527 0.778 49.6  99.3 11.5
9 8 17.08 +34 31.0 0.550 0.770 49.2  98.1 11.6
11 8 24.20 +32 35.6 0.573 0.763 48.9  96.7 11.6
13 8 30.95 +30 45.0 0.597 0.758 48.6  95.2 11.7
15 8 37.40 +28 59.1 0.621 0.755 48.5  93.4 11.7

1994 August 30 Daniel W. E. Green